# Differentially Private Federated Learning of Diffusion Models for Synthetic Tabular Data Generation

### v1

@@ 4.1. Datasets and Data Preparation
    Census Income Data [^4] https://archive.ics.uci.edu/dataset/2/adult ($\mathcal{D}_{B}$): This dataset contains demographic information from the 1994 U.S. Census to predict whether a person earns more than \$50,000 per year. In total, there are 32,561 records each encompassing 10 categorical and 3 numerical attributes.
  3. (3)
@@
    Philadelphia City Payments Data [^5] https://tinyurl.com/bdz2xdbx ($\mathcal{D}_{C}$): This dataset consists of 238,894 payment records. The payments were generated by 58 distinct city departments in 2017. Each payment includes 10 categorical and 1 numerical attribute(s).
  4. (4)
@@
    Marketing Data [^6] https://tinyurl.com/zx4u8tf5 ($\mathcal{D}_{D}$) This dataset contains 45,211 customer records of a bank from 2008 to 2010, used to predict whether a client will subscribe to a term deposit. Each record includes 10 categorical and 6 numerical attributes.
@@
Figure 3. Non-IID data distribution among (3, 5, and 10) client settings. Each bar indicates the percentage of data allocated to each client $\omega_{i}$ across four datasets.
@@
To simulate a realistic non-IID and unbalanced data environment for federated training, each dataset is partitioned based on a categorical feature. The descriptive statistics on the non-IID data partitioning schemes of 3, 5, and 10 client settings are presented in Figure 3. To standardize the numeric attributes, we employed quantile transformations, as implemented in the scikit-learn library.  [^7] https://tinyurl.com/ht9pz8m5 For the categorical attributes, we utilized embedding techniques following the approach outlined by Sattarov et al. (2023).
4.2. Model Architecture and Hyperparameters
@@
In the following, we detail the architecture and the specific hyperparameters chosen in DP-Fed-FinDiff model optimization.
@@
Diffusion Model. [^8] Model parameter optimization is conducted using PyTorch v2.2.1 (Paszke et al., 2019) The architecture for all datasets consists of three layers, each comprising 1024 neurons, except for $\mathcal{D}_{C}$, which contains 2048 neurons. The models are trained for up to $R=3,000$ communication rounds utilizing a mini-batch size of 16. The Adam optimizer (Kingma and Ba, 2015) is utilized with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ The hyperparameters for the diffusion model FinDiff model are adopted from (Sattarov et al., 2023). These settings include 500 diffusion steps ( $T=500$ ) and a linear learning-rate scheduler with initial and final rates of $\beta_{start}=0.0001$ and $\beta_{end}=0.02$ , respectively. Each categorical attribute is represented as a 2-dimensional embedding.
@@
Federated Learning. [^9] The federated learning scenario is simulated using the Flower framework v1.7.0 (Beutel et al., 2022).  [^10] https://flower.ai/docs/framework/explanation-differential-privacy.html In each communication round $r=1,\ldots,R$ , each client $\omega_{i}$ performs $\gamma=1,\dots,\Gamma$ local optimization updates on its model $\theta^{\omega}_{i}$ before sharing the updated parameters. The number of client optimization updates is evaluated across various settings $\Gamma\in[10,50,100,500,1000]$ Configurations with different numbers of clients $\lambda\in[3,5,10]$ are also examined. Four distinct federated optimization strategies are explored: Federated Averaging (FedAvg(McMahan et al., 2017), Federated Adam (FedAdam(Reddi et al., 2020), Federated Proximal (FedProx(Li et al., 2020), and Federated Yogi (FedYogi(Reddi et al., 2020). FedAvg aggregates client models by computing the weighted average of their parameters. FedAdam, an extension of the Adam optimizer, utilizes default hyperparameters $\beta_{1}=0.9$ , $\beta_{2}=0.999$ , and $\epsilon=1e-8$ FedProx introduces a proximal term to address client heterogeneity, employing a default $\mu=0.01$ Lastly, FedYogi adapts the Yogi optimizer using default parameters $\beta_{1}=0.9$ , $\beta_{2}=0.999$ , and $\epsilon=1e-8$
@@
Differential Privacy. [^11] For training and accounting of the differential privacy we use Opacus v1.4.1 (Yousefpour et al., 2021). We adopt the privacy settings from (Dockhorn et al., 2022), training models with $\varepsilon$ values of $\{0.2,1,10\}$, corresponding to high, moderate, and low privacy levels. The probability of information leakage is set to the reciprocal of the number of training samples, $\delta=N^{-1}$, a common heuristic in practice.
@@
Figure 4. Comparative evaluation of local optimization updates $\Gamma\in[10,50,100,500,1000]$ between (a) non-IID and (b) IID settings, highlighting their impact on privacy, utility, and fidelity. In IID settings, increasing the number of local updates consistently improves performance, whereas, in non-IID settings, performance degrades after reaching a certain threshold.
Figure 5. Heatmaps illustrating the impact of Differential Privacy (DP) budgets $\varepsilon\in[0.2,1,10]$ and $\infty$ (no DP) on various datasets across three key metrics: Privacy, Utility, and Fidelity in non-IID settings. It is observed that as the DP budget decreases (i.e., as $\varepsilon$ values lower from $\infty$ to 0.2), privacy protection improves, while fidelity and utility decline.
@@
Figure 6. t-SNE visualization of (a) original Credit Default data ($\mathcal{D}_{A}$), (b) synthetic data generated without differential privacy (DP), and (c, d, e) synthetic data generated with DP using federated optimization across 5 clients ($\lambda$=5). As privacy levels increase (reflected by a decrease in $\varepsilon$), the structural integrity of the synthetic data diminishes, resulting from the increased DP noise.
4.3. Evaluation Metrics
A comprehensive set of standard evaluation metrics, including privacy, utility, and fidelity, is employed to evaluate the model’s effectiveness. These metrics represent diverse aspects of data generation quality, providing a holistic view of model performance.
Privacy. [^12] The estimation of privacy risks was conducted using the Anonymeter library (Giomi et al., 2023). The privacy metric quantifies the extent to which synthetic data prevents the identification of original data entries. In this study, privacy is assessed using three key indicators of factual anonymization as outlined by the GDPR (Article 29 Data Protection Working Party, 2014). Specifically, we employ privacy evaluators to measure the risks of (i) singling out, (ii) linkability, and (iii) inference that could potentially affect data donors following the release of synthetic datasets. These risk measurements are defined as follows, where $s$ denotes the synthetic dataset and $x$ represents the real dataset:
@@ 4.3. Evaluation Metrics
  • Linkability Risk measures the proportion of successful attribute linkages $LR(x,s)$ between synthetic and real records. This risk evaluates the potential for linking a synthetic record to a real one by matching shared attributes, using a subset of 10 attributes (6 for dataset $\mathcal{D_{C}}$).
@@
For each evaluator, the risk is estimated by performing 500 attacks on each record, and the overall privacy risk is computed as the mean score across all synthetic data points. The comprehensive privacy score is the aggregated risk from all three evaluators:
(7) $\Pi=\frac{1}{3}(SOR(x,s)+LR(x,s)+IR(x,s)).$
@@
Utility. The effectiveness of synthetic data is determined by its utility, a measure of how functionally equivalent it is to real-world data. This utility is quantified by training machine learning models on synthetic datasets and then assessing their performance on original datasets. In this study, we operationalize utility as the performance of classifiers trained on synthetic data ($S_{Train}$) that shares dimensional consistency with the real training set and then evaluated against the actual test set ($X_{Test}$). This process evaluates the synthetic data’s efficacy in replicating the statistical properties necessary for accurate model training. The average accuracy across all classifiers is computed to represent the overall utility of the synthetic data, formalized as:
(8) $\Phi=\frac{1}{N}\sum_{i=1}^{N}\Theta_{i}(S_{Train},X_{Test}).$
@@
Here, $\Phi$ represents the utility score, and $\Theta_{i}$ denotes the accuracy of the $i$-th classifier. To provide a comprehensive evaluation, we selected $N$=5 classifiers for this study, namely Random Forest, Decision Trees, Logistic Regression, Ada Boost, and MLP Classifier.
@@
Fidelity. [^13] The row fidelity computation was performed using Dython library v0.7.5 (Zychlinski, 2018). Fidelity assesses how closely synthetic data emulates real data, considering both column-level and row-level comparisons. For column fidelity, the similarity between corresponding columns in synthetic and real datasets is evaluated. Numeric attributes employ the Wasserstein similarity, represented as $WS(x^{d},s^{d})$ , to measure the distance between the distributions of numeric attributes. The Jensen-Shannon divergence, denoted as $JS(x^{d},s^{d})$ quantifies differences in categorical attributes. These metrics were combined to form the column fidelity score $\Omega_{col}$ as:
(9) $\Omega_{col}=\begin{cases}1-WS(x^{d},s^{d})&\text{if $d$ is num.}\\ 1-JS(x^{d},s^{d})&\text{if $d$ is cat.}\end{cases}$
@@
The overall fidelity for columns in synthetic dataset $S$ is the mean of $\Omega_{col}(x^{d},s^{d})$ across all attributes. Row fidelity focuses on correlations between column pairs. For numeric attributes, the Pearson Correlation between pairs $\rho(x^{a},x^{b})$ is used. The discrepancy in correlations for real and synthetic pairs, $PC(x^{a,b},s^{a,b})=|\rho(x^{a},x^{b})-\rho(s^{a},s^{b})|$ , quantifies this aspect. The Theil U coefficient, also known as Theil’s uncertainty coefficient, similarly quantifies the association between two categorical variables, denoted as $TU(x^{a,b},s^{a,b}$) . Similarly to the column fidelity, these metrics are combined to form the row fidelity:
(10) $\Omega_{row}=\begin{cases}1-PC(x^{a,b},s^{a,b})&\text{if $d$ is num.}\\ 1-TU(x^{a,b},s^{a,b})&\text{if $d$ is cat.}\end{cases}$
@@
The total row fidelity for the dataset $S$ is the average of $\Omega_{row}(x^{a,b},s^{a,b})$ across all attribute pairs. Finally, the aggregate fidelity score, denoted as $\Omega(X,S)$ , is the mean of column and row fidelity.
5. Experimental results.
@@
This section presents the results of the experiments, demonstrating the efficacy of the DP-Fed-FinDiff model and providing quantitative analyses. The conducted experiments are accompanied by three Research Questions (RQ) described below.
@@
Figure 7. Singling Out, Linkability, and Inference risk evaluation under various privacy budgets ($\varepsilon$$\infty$ denotes no DP) across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$). It is observed that the privacy risk metrics increase as the privacy budget increases.
@@
RQ 1: How does varying the number of local optimization updates impact the training of diffusion models for tabular data in a federated learning setup employing differential privacy?
@@
For utility and fidelity, increasing the number of client updates in the IID setting (see 4(b)) consistently improved performance. Each client received an IID data partition, which helped maintain model quality. However, in the non-IID setting (see 4(a)), performance declined after a threshold of $\Gamma=100$. The decline is caused by heterogeneous data distributions among clients. As a result, the clients drift away from a globally optimal model (Karimireddy et al., 2019), leading to unstable and slow convergence.
Our findings indicate that 100 local updates provide an optimal balance between privacy, utility, and fidelity in both IID and non-IID settings. Additionally, this choice significantly reduced the overall training time from 28 hours ($\Gamma=1000$) to 3.5 hours ($\Gamma=100$).
@@
We assessed the impact of differential privacy on fidelity, utility, and privacy risks across three privacy budgets $\varepsilon\in[0.2,1,10]$ and without DP ($\varepsilon=\infty$) in a non-IID setting. The number of federated clients was fixed at $\lambda=5$ with client optimization rounds set to $\Gamma=100$. Privacy risks were measured by estimating Singling Out, Linkability, and Inference risks.
Figure 8. Privacy, utility, and fidelity evaluation using a varying number of federated clients $\lambda\in[3,5,10]$ across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$) and non-IID setting. We kept the privacy budget $\epsilon=1$, local updates $\Gamma=100$ and FedAvg strategy fixed.
Figure 9. Privacy, utility, and fidelity evaluation using different federated strategies across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$) and non-IID setting. We kept the privacy budget $\epsilon=1$, local optimization updates $\Gamma=100$ and number of clients $\lambda=5$ fixed.
@@
Additionally, lower DP budgets ($\varepsilon=0.2$ and $\varepsilon=1.0$) mitigate Singling Out, Linkability, and Inference risks, demonstrating the need for stringent DP constraints to minimize re-identification, linkage, and inference threats (see Figure 7). Higher DP budgets ($\varepsilon=10$ and $\infty$) reduce noise, enhancing data utility but compromising privacy. These findings highlight the trade-off between privacy and utility in synthetic data generation, emphasizing the importance of careful DP parameter selection.
@@
Additionally, fidelity and utility scores also show improvement with a higher number of clients. We attribute this to the regularizing effect of differential privacy, which reduces the client drift effect and results in more consistent local models. However, these benefits have limits. When expanding from five to ten clients, the gains in utility become marginal and can even decline. This performance degradation is likely due to the increased complexity of aggregating updates from a larger number of clients.
@@
In summary, the evaluation of RQ 1, 2, and 3 reveals the trade-offs in federated learning with differential privacy, highlighting the critical balance required between privacy and data quality, as well as the influence of optimization strategies and federated configurations on the overall DP-Fed-FinDiff model performance.
6. Conclusion
This study introduces the DP-Fed-FinDiff framework, a novel integration of (i) Differential Privacy (DP), (ii) Federated Learning (FL), and (iii) Denoising Diffusion Probabilistic Models (DDPMs) for generating high-fidelity synthetic tabular data. The framework addresses the critical need for privacy-preserving data generation in finance and other sensitive but high-stake domains.
Our comprehensive evaluations revealed the trade-offs between data quality and privacy. The model proves to be a robust solution for generating privacy-preserving synthetic data, paving the way for secure data sharing and advanced analytics in high-stakes environments. Future work will explore adaptive strategies to dynamically balance privacy and data quality, further enhancing the applicability of federated learning in diverse settings.

### v2

@@ IV-A Datasets and Data Preparation
    Census Income [^4] https://archive.ics.uci.edu/dataset/2/adult ($\mathcal{D}_{B}$): This dataset contains demographic information from the 1994 U.S. Census to predict whether a person earns more than \$50,000 per year. In total, there are 32,561 records each encompassing 10 categorical and 3 numerical attributes.
  3. 3.
@@
    Philadelphia City Payments [^5] https://tinyurl.com/bdz2xdbx ($\mathcal{D}_{C}$): This dataset consists of 238,894 records. The payments were generated by 58 distinct city departments in 2017. Each payment includes 10 categorical and 1 numerical attribute(s).
  4. 4.
@@
    Marketing [^6] https://tinyurl.com/zx4u8tf5 ($\mathcal{D}_{D}$) This dataset contains 45,211 customer records of a bank from 2008 to 2010, used to predict whether a client will subscribe to a term deposit. Each record includes 10 categorical and 6 numerical attributes.
@@
Figure 3: Non-IID data distribution among (3, 5, and 10) client settings. Each bar indicates the percentage of data allocated to each client $\omega_{i}$ across four datasets.
@@
To simulate a realistic non-IID and unbalanced data environment for federated training, each dataset is partitioned based on a categorical feature. The descriptive statistics on the non-IID data partitioning schemes of 3, 5, and 10 client settings are presented in Figure 3. To standardize the numeric attributes, we employed quantile transformations [^7] https://tinyurl.com/ht9pz8m5. For embedding the categorical attributes, we followed the same approach as outlined in [13].
IV-B Model Architecture and Hyperparameters
@@
Next, we detail the architecture and hyperparameters chosen in DP-FedTabDiff model optimization.
@@
Diffusion Model. [^8] Model parameter optimization is conducted using PyTorch v2.2.1 [46]. The architecture for all datasets consists of three layers, each comprising 1024 neurons, except for $\mathcal{D}_{C}$, which contains 2048 neurons. The models are trained for up to $R=3,000$ communication rounds utilizing a mini-batch size of 16. The Adam optimizer [47] is utilized with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ . The hyperparameters of the underlying diffusion model FinDiff are adopted from [13]. These settings include 500 diffusion steps ( $T=500$ ) and a linear learning-rate scheduler with initial and final rates of $\beta_{start}=0.0001$ and $\beta_{end}=0.02$ , respectively. Each categorical attribute is represented as a 2-dimensional embedding.
@@
Federated Learning. [^9] Federated learning is simulated using the Flower framework v1.7.0 [48]. In each communication round $r=1,\ldots,R$ , a random client $\omega_{i}$ performs $\gamma=1,\dots,\Gamma$ local optimization updates on its model $\theta^{\omega}_{i}$ before sharing the updated parameters. The number of client optimization updates is evaluated across various settings $\Gamma\in[10,50,100,500,1000]$ . Configurations with different numbers of clients $\lambda\in[3,5,10]$ are also examined. Four distinct federated optimization strategies are explored: Federated Averaging (FedAvg[5], Federated Adam (FedAdam[49], Federated Proximal (FedProx[50], and Federated Yogi (FedYogi[49]. FedAvg aggregates client models by computing the weighted average of their parameters. FedAdam, an extension of the Adam optimizer, utilizes default hyperparameters $\beta_{1}=0.9$ , $\beta_{2}=0.999$ , and $\epsilon=1e-8$ . FedProx introduces a proximal term to address client heterogeneity, employing a default $\mu=0.01$ . Lastly, FedYogi adapts the Yogi optimizer using default parameters $\beta_{1}=0.9$ , $\beta_{2}=0.999$ , and $\epsilon=1e-8$ .
@@
Differential Privacy. [^10] Training with differential privacy is performed using Opacus v1.4.1 [51]. We adopt the privacy settings from [34], training models with $\varepsilon\in\{0.2,1,10\}$, corresponding to high, moderate, and low privacy levels. The probability of information leakage is set to the reciprocal of the number of training samples, $\delta=N^{-1}$, a common heuristic in practice.
@@
Figure 4: Comparative evaluation of local optimization updates $\Gamma\in[10,50,100,500,1000]$ between (a) non-IID and (b) IID settings, highlighting their impact on privacy, utility, and fidelity. In IID settings, increasing the number of local updates consistently improves performance, whereas, in non-IID settings, performance degrades after reaching a certain threshold.
Figure 5: Heatmaps illustrating the impact of DP budgets $\varepsilon\in[0.2,1,10]$ and $\infty$ (no DP) on various datasets across Privacy, Utility, and Fidelity in non-IID settings. It is observed that as the DP budget decreases (i.e. lower $\varepsilon$ values), privacy protection improves, while fidelity and utility decline.
@@
Figure 6: t-SNE visualization of (a) dataset $\mathcal{D}_{A}$, (b) synthetic data without DP, and (c-e) with DP using federated optimization across 5 clients ($\lambda$=5). As privacy increases (i.e., lower $\varepsilon$), DP noise degrades the structural integrity of the synthetic data.
IV-C Evaluation Metrics
A comprehensive set of evaluation metrics, including privacy, utility, and fidelity, is used to assess the model’s effectiveness, offering a holistic view of data generation quality.
Privacy. [^11] The privacy risks estimation was conducted using the Anonymeter [52]. The privacy metric quantifies the extent to which synthetic data prevents the identification of original data entries. In this study, privacy is assessed using three key indicators of factual anonymization as outlined by the GDPR [53]. Specifically, we employ privacy evaluators to measure the risks of (i) singling out, (ii) linkability, and (iii) inference that could potentially affect data donors following the release of synthetic datasets. These risk measures are defined as follows, with $\bm{S}$ for synthetic dataset and $\bm{X}$ for real dataset:
@@ IV-C Evaluation Metrics
  • Linkability Risk measures the proportion of successful attribute linkages $LR(X,S)$ between synthetic and real records. It reflects the potential to link synthetic to real records via shared attributes, using a subset of 10 attributes (6 for the dataset $\mathcal{D_{C}}$).
@@
Each evaluator performs 500 attacks per record, computing the mean risk over all synthetic samples. The final privacy score aggregates results from all three evaluators, as defined by:
$\Pi=\frac{1}{3}(SOR(X,S)+LR(X,S)+IR(X,S)).$ (7)
@@
Utility. The effectiveness of synthetic data is determined by its utility, a measure of how functionally equivalent it is to real-world data. We quantify it by training classifiers on synthetic data ($S_{\text{train}}$), aligned dimensionally with the real training set, and evaluating on the real test set ($X_{\text{test}}$). This assesses how well statistical properties are preserved for model training. The average classifier accuracy represents overall utility, formalized as:
$\Phi=\frac{1}{N}\sum_{i=1}^{N}\Theta_{i}(S_{\text{train}},X_{\text{test}}).$ (8)
@@
Here, $\Phi$ represents the utility score, and $\Theta_{i}$ denotes the accuracy of the $i$-th classifier. To provide a comprehensive evaluation, we selected NN=5 classifiers for this study, namely Random Forest, Decision Trees, Logistic Regression, Ada Boost, and MLP Classifier.
@@
Fidelity. [^12] The row fidelity was computed using Dython library v0.7.5 [54]. Fidelity assesses how closely synthetic data emulates real data, considering both column-level and row-level comparisons. For column fidelity, the similarity between corresponding columns in synthetic and real datasets is evaluated. To measure the distance between the distributions of numeric attributes the Wasserstein similarity $WS(x^{d},s^{d})$ is employed. The Jensen-Shannon divergence, denoted as $JS(x^{d},s^{d})$ quantifies differences in categorical attributes. These metrics were combined to form the column fidelity score $\Omega_{col}$ as:
$\Omega_{col}=\begin{cases}1-WS(x^{d},s^{d})&\text{if $d$ is num.}\\ 1-JS(x^{d},s^{d})&\text{if $d$ is cat.}\end{cases}$ (9)
@@
The overall fidelity for columns in synthetic dataset $S$ is the mean of $\Omega_{col}(x^{d},s^{d})$ across all attributes. Row fidelity focuses on correlations between column pairs. For numeric attributes, the Pearson Correlation between pairs $\rho(x^{a},x^{b})$ is used. The discrepancy in correlations for real and synthetic pairs, $PC(x^{a,b},s^{a,b})=|\rho(x^{a},x^{b})-\rho(s^{a},s^{b})|$ , quantifies this aspect. The Theil U coefficient quantifies the association between two categorical variables, denoted as $TU(x^{a,b},s^{a,b}$) .
$\Omega_{row}=\begin{cases}1-PC(x^{a,b},s^{a,b})&\text{if $d$ is num.}\\ 1-TU(x^{a,b},s^{a,b})&\text{if $d$ is cat.}\end{cases}$ (10)
@@
The total row fidelity for the dataset $S$ is the average of $\Omega_{row}(x^{a,b},s^{a,b})$ across all attribute pairs. Finally, the aggregate fidelity score $\Omega(X,S)$ is the mean of column and row fidelity.
V Experimental results.
@@
This section presents the results of the experiments, demonstrating the efficacy of the DP-FedTabDiff model and providing quantitative analyses. The conducted experiments are accompanied by three Research Questions (RQ).
@@
Figure 7: Singling Out, Linkability, and Inference risk evaluation under various privacy budgets ($\varepsilon$$\infty$ denotes no DP) across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$). It is observed that the privacy risk metrics increase as the privacy budget increases.
@@
RQ 1: How does varying number of local optimization updates impact the training of diffusion models for tabular data in a federated learning setup employing differential privacy?
@@
For utility and fidelity, increasing the number of client updates in the IID setting (see 4(b)) consistently improved performance. Each client received an IID data partition, which helped maintain model quality. However, in the non-IID setting (see 4(a)), performance declined after a threshold of $\Gamma=100$. The decline is caused by heterogeneous data distributions among clients. As a result, the clients drift away from a globally optimal model [55], leading to unstable and slow convergence.
Our findings indicate that 100 local updates provide an optimal trade-off between privacy, utility, and fidelity in both IID and non-IID settings. Additionally, this choice significantly reduced the overall training time from 28 hours ($\Gamma=1000$) to 3.5 hours ($\Gamma=100$).
@@
We assessed the impact of DP on fidelity, utility, and privacy risks across three privacy budgets $\varepsilon\in[0.2,1,10]$ and without DP ($\varepsilon=\infty$) in a non-IID setting. The number of federated clients was fixed at $\lambda=5$ with client optimization rounds set to $\Gamma=100$. Privacy risks were measured by estimating Singling Out, Linkability, and Inference risks.
Figure 8: Privacy, utility, and fidelity evaluation using a varying number of federated clients $\lambda\in[3,5,10]$ across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$) and non-IID setting. We kept the privacy budget $\epsilon=1$, local updates $\Gamma=100$ and FedAvg strategy fixed.
Figure 9: Privacy, utility, and fidelity evaluation using different federated strategies across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$) and non-IID setting. We kept the privacy budget $\epsilon=1$, local optimization updates $\Gamma=100$ and number of clients $\lambda=5$ fixed.
@@
Additionally, lower DP budgets ($\varepsilon=0.2$ and $\varepsilon=1.0$) mitigate Singling Out, Linkability, and Inference risks, demonstrating the need for stringent DP constraints to minimize re-identification, linkage, and inference threats (see Figure 7). Higher DP budgets ($\varepsilon=10$ and $\infty$) reduce noise, enhancing data utility but compromising privacy. These results underscore the privacy-utility trade-off in synthetic data generation and the need for careful DP parameter tuning.
@@
Additionally, fidelity and utility scores improve with a higher number of clients. We attribute this to the regularizing effect of DP, which reduces the client drift effect and results in more consistent local models. However, these benefits have limits. When expanding from five to ten clients, the gains in utility become marginal and can even decline. This performance degradation is likely due to the increased complexity of aggregating updates from a larger number of clients.
@@
In summary, the evaluation of RQ 1, 2, and 3 reveals the trade-offs in federated learning with DP, highlighting the critical balance between privacy and data quality, as well as the influence of optimization strategies and federated configurations on the overall DP-FedTabDiff model performance.
VI Conclusion
We introduced DP-FedTabDiff, a novel framework that unifies Differential Privacy (DP), Federated Learning (FL), and Denoising Diffusion Probabilistic Models (DDPMs) to enable high-fidelity and privacy-preserving synthetic tabular data generation. By jointly addressing data utility and privacy, our framework offers a practical solution for secure data sharing and downstream analytics in sensitive, high-stakes domains such as finance or healthcare.
Our comprehensive evaluations revealed the trade-offs between data quality and privacy, showing that DP-FedTabDiff maintains strong performance under realistic privacy budgets. These results highlight its robustness and potential as a foundation for privacy-aware generative modeling in distributed settings. Future work will explore adaptive strategies to dynamically balance privacy and data quality, further enhancing the applicability of federated learning in diverse settings.